OpenStudy (anonymous):

i have a right angle with a line through it one angle equals (3x) the other equals (2x) so 90 =3x+2x?

OpenStudy (anonymous):

yes, if i read your question correctly.

OpenStudy (anonymous):

hi thats correct,..3x+2x=90... 5x=90 x=90/5....calculator or using radian measure 3x+2x=2pi 5x=2pi x=2pi/5

OpenStudy (anonymous):

hi thats correct,..3x+2x=90... 5x=90 x=90/5....calculator or using radian measure 3x+2x=2pi 5x=2pi x=2pi/5

OpenStudy (anonymous):

hi thats correct,..3x+2x=90... 5x=90 x=90/5....calculator or using radian measure 3x+2x=2pi 5x=2pi x=2pi/5